# Gaussian Process Upper Confidence Bound Achieves Nearly-Optimal Regret in Noise-Free Gaussian Process Bandits

### v1

@@ Abstract
We study the noise-free Gaussian Process (GP) bandits problem, in which the learner seeks to minimize regret through noise-free observations of the black-box objective function lying on the known reproducing kernel Hilbert space (RKHS). Gaussian process upper confidence bound (GP-UCB) is the well-known GP-bandits algorithm whose query points are adaptively chosen based on the GP-based upper confidence bound score. Although several existing works have reported the practical success of GP-UCB, the current theoretical results indicate its suboptimal performance. However, GP-UCB tends to perform well empirically compared with other nearly optimal noise-free algorithms that rely on a non-adaptive sampling scheme of query points. This paper resolves this gap between theoretical and empirical performance by showing the nearly optimal regret upper bound of noise-free GP-UCB. Specifically, our analysis shows the first constant cumulative regret in the noise-free settings for the squared exponential kernel and Matérn kernel with some degree of smoothness.
@@ 1 Introduction
This paper studies the noise-free Gaussian Process (GP) bandits problem, where the learner seeks to minimize regret through noise-free observations of the black-box objective function. Several existing works tackle this problem, and some of them (Salgia et al., 2024; Iwazaki and Takeno, 2025) show the algorithms whose regret nearly matches the conjectured lower bound of (Vakili, 2022). For ease of theoretical analysis, these algorithms rely on the non-adaptive sampling scheme, whose query points are chosen independently from the observed function values, such as the uniform sampling (Salgia et al., 2024) or maximum variance reduction (Iwazaki and Takeno, 2025). Although the theoretical superiority of such non-adaptive algorithms is shown, their empirical performances are known to be worse than those of adaptive strategy, whose current regret upper bound may have significant room for improvement. From this motivation, our work aims to show the nearly optimal regret of GP upper confidence bound (GP-UCB) (Srinivas et al., 2010), which is one of the well-known adaptive GP bandit algorithms, and its existing guarantees only show strictly sub-optimal regret in noise-free setting (Lyu et al., 2019; Kim and Sanz-Alonso, 2024).
@@
  • We give a refined regret analysis of GP-UCB (Theorems 1 and 2), which matches both the conjectured cumulative regret lower bounds of (Vakili, 2022) and the simple regret lower bound of (Bull, 2011) up to polylogarithmic factors in the Matérn kernel. Regarding cumulative regret, our analysis shows that GP-UCB achieves the constant $O(1)$ regret under squared exponential and Matérn kernel with $d>\nu$. Here, $d$ and $\nu$ denote the dimension and smoothness parameter of the Matérn kernel, respectively. The results are summarized in Tables 1 and 2.
  • Specifically, our key theoretical contribution is the new algorithm-independent upper bound for the observed posterior standard deviations (Lemmas 35). As discussed in Remark 1, this result has the potential to translate existing confidence bound-based algorithms for noisy settings into nearly optimal noise-free variants beyond the analysis of GP-UCB.
  • We also show the first algorithm-independent lower bound for deterministic regret under Matérn kernel with $d>\nu$ (Theorem 7). This result formally validates the conjectured lower bound of (Vakili, 2022) up to the poly-logarithmic factor and the near-optimality of our analysis of GP-UCB.
@@
A lot of existing works study the theory for the noisy GP bandits (Srinivas et al., 2010; Valko et al., 2013; Chowdhury and Gopalan, 2017; Scarlett et al., 2017; Li and Scarlett, 2022). Regarding noise-free settings, to our knowledge, (Bull, 2011) is the first work that shows both the upper bound and the lower bound for simple regret via the expected improvement (EI) strategy. After that, the analysis of the cumulative regret is shown in (Lyu et al., 2019) with GP-UCB. Recently, several works studied the improved algorithm to achieve superior regret to the result of (Lyu et al., 2019) based on the conjectured lower bound provided in (Vakili, 2022). Although several works show the conjectured nearly-optimal algorithm (Salgia et al., 2024; Iwazaki and Takeno, 2025), their algorithms are based on a non-adaptive sampling scheme, whose inferior performances were reported in the existing work (Li and Scarlett, 2022). On the other hand, our proof technique relates to the analysis of (Iwazaki and Takeno, 2025) for the non-adaptive maximum variance reduction algorithm. The theoretical analysis of this paper can be interpreted as a generalization of the result of (Iwazaki and Takeno, 2025) tailored to a noise-free setting. Finally, in contrast to the frequentist assumption that this paper focuses on, some existing works also study the Bayesian assumption of the objective function (Grünewälder et al., 2010; Srinivas et al., 2010; De Freitas et al., 2012; Scarlett, 2018).
Table 1: Comparison between existing noiseless algorithms’ guarantees for cumulative regret and our result. In all algorithms, the smoothness parameter of the Matérn kernel is assumed to be $\nu>1/2$. Furthermore, $d$, $\ell$, $\nu$, and $B$ are supposed to be $\Theta(1)$ here. “Type” column shows that the regret guarantee is (D)eterministic or (P)robabilistic. Here, we describe the regret bound as “deterministic” if the regret upper or lower bound always holds without probabilistic arguments. Throughout this paper, $\widetilde{O}(\cdot)$ denotes the order notation whose poly-logarithmic dependence is ignored.
Algorithm Regret (SE) Regret (Matérn) Type Remark
@@
(Vakili, 2022)
Kernel-AMM-UCB $O\left(\ln^{d+1}T\right)$ $\widetilde{O}\left(T^{\frac{\nu d+d^{2}}{2\nu^{2}+2\nu d+d^{2}}}\right)$ D
(Flynn and Reeb, 2024)
REDS N/A $\widetilde{O}\left(T^{\frac{d-\nu}{d}}\right)$ $O\left(\ln^{\frac{5}{2}}T\right)$ $O\left(\ln^{\frac{3}{2}}T\right)$ P Assumption for
(Salgia et al., 2024) level-set is required.
PE $O\left(\ln T\right)$ $\widetilde{O}\left(T^{\frac{d-\nu}{d}}\right)$ $O\left(\ln^{2+\alpha}T\right)$ $O\left(\ln T\right)$ D $\alpha>0$ is an arbitrarily
(Iwazaki and Takeno, 2025) fixed constant.
@@
(Vakili, 2022)
Lower Bound N/A $\Omega\left(T^{\frac{d-\nu}{d}}\right)$ N/A N/A D
Table 2: Comparison between existing noiseless algorithms’ guarantees for simple regret and our result. In all algorithms except for GP-UCB+ and EXPLOIT+, the smoothness parameter of the Matérn kernel is assumed to be $\nu>1/2$.
Algorithm Regret (SE) Regret (Matérn) Type Remark
@@
(Lyu et al., 2019)
(Kim and Sanz-Alonso, 2024)
Kernel-AMM-UCB $O\left(\frac{\ln^{d+1}T}{T}\right)$ $\widetilde{O}\left(T^{-\frac{\nu d+2\nu^{2}}{2\nu^{2}+2\nu d+d^{2}}}\right)$ D
(Flynn and Reeb, 2024)
GP-UCB+, $O\left(\exp\left(-CT^{\frac{1}{d}-\alpha}\right)\right)$ $O\left(T^{-\frac{\nu}{d}+\alpha}\right)$ P $\alpha>0$ is an arbitrarily
EXPLOIT+ fixed constant.
(Kim and Sanz-Alonso, 2024) $C>0$ is some constant.
MVR $O\left(\exp\left(-\frac{1}{2}T^{\frac{1}{d+1}}\ln^{-\alpha}T\right)\right)$ $\widetilde{O}\left(T^{-\frac{\nu}{d}}\right)$ D $\alpha>0$ is an arbitrarily
(Iwazaki and Takeno, 2025) fixed constant.
GP-UCB $O\left(\sqrt{T}\exp\left(-\frac{1}{2}\widetilde{C}_{\mathrm{SE}}T^{\frac{1}{d+1}}\right)\right)$ $\widetilde{O}\left(T^{-\frac{\nu}{d}}\right)$ D $\widetilde{C}_{\mathrm{SE}}>0$ is defined
(Our analysis) in Theorem 2.
@@
(Bull, 2011)
@@ 2 Preliminaries
This paper studies the noise-free Gaussian process (GP) bandit problem. Let $f:\mathcal{X}\rightarrow\mathbb{R}$ be a black-box objective function whose input domain $\mathcal{X}\subset\mathbb{R}^{d}$ is compact. At each step $t\in\mathbb{N}_{+}$, the learner chooses query point $\bm{x}_{t}\in\mathcal{X}$; after that, the corresponding function value $f(\bm{x}_{t})$ is returned. Under the total step size $T\in\mathbb{N}_{+}$, the learner’s goal is to minimize one of the following metrics: cumulative regret $R_{T}$ and simple regret $r_{T}$, which are respectively defined as
$R_{T}$ $=\sum_{t\in[T]}f(\bm{x}^{\ast})-f(\bm{x}_{t}),$ (1)
@@
Here, we define $[T]\coloneqq\{1,\ldots,T\}$ and $\bm{x}^{\ast}\in\mathrm{arg~{}max}_{\bm{x}\in\mathcal{X}}f(\bm{x})$. Furthermore, $\widehat{\bm{x}}_{T}\in\mathcal{X}$ is the estimated maximizer returned by the algorithm at the end of step $T$.
Regularity Assumptions.
We suppose the following assumption for the underlying objective function.
Assumption 1.
The objective function $f$ is an element of reproducing kernel Hilbert space (RKHS), endowed with a known positive definite kernel $k:\mathcal{X}\times\mathcal{X}\rightarrow\mathbb{R}$. Furthermore, assume $k(\bm{x},\bm{x})\leq 1$ for all $\bm{x}\in\mathcal{X}$, and $\|f\|_{k}\leq B<\infty$, where $\|f\|_{k}$ denote the RKHS norm of $f$.
Assumption 1 is a standard assumption in GP-bandits literature (Srinivas et al., 2010; Scarlett et al., 2017; Chowdhury and Gopalan, 2017). Specifically, we focus on the following squared exponential (SE) kernel $k_{\mathrm{SE}}$ and Matérn kernel $k_{\mathrm{Mat\acute{e}rn}}$:
$k_{\mathrm{SE}}(\bm{x},\widetilde{\bm{x}})$ $=\exp\left(\frac{\|\bm{x}-\widetilde{\bm{x}}\|_{2}^{2}}{2\ell^{2}}\right),$ (3)
$k_{\mathrm{Mat\acute{e}rn}}(\bm{x},\widetilde{\bm{x}})$ $=\frac{2^{1-\nu}}{\Gamma(\nu)}\left(\frac{\sqrt{2\nu}\|\bm{x}-\widetilde{\bm{x}}\|_{2}}{\ell}\right)J_{\nu}\left(\frac{\sqrt{2\nu}\|\bm{x}-\widetilde{\bm{x}}\|_{2}}{\ell}\right),$ (4)
where $\ell>0$ and $\nu>0$ are the lengthscale and smoothness parameters, respectively. Futhermore, $J_{\nu}(\cdot)$ and $\Gamma(\cdot)$ respectively denote modified Bessel and Gamma function.
@@
GP model is a useful Bayesian model for quantifying both prediction and uncertainty of underlying data. It is often leveraged to construct an algorithm in existing GP-bandits works. Let us assume the Bayesian assumption of $f$ that $f$ follows mean-zero GP, characterized by the kernel function $k$. Then, given the input $\mathbf{X}_{t}=(\bm{x}_{1},\ldots,\bm{x}_{t})$ and corresponding outputs, the posterior distribution of $f$ is again GP, whose mean and variance function of $f(\bm{x})$ are defined as
@@
where $\bm{k}(\bm{x},\mathcal{E}(\mathbf{X}_{t}))\coloneqq[k(\widetilde{\bm{x}},\bm{x})]_{\widetilde{\bm{x}}\in\mathcal{E}(\mathbf{X}_{t})}$, $\bm{f}(\mathcal{E}(\mathbf{X}_{t}))=[f(\widetilde{\bm{x}})]_{\widetilde{\bm{x}}\in\mathcal{E}(\mathbf{X}_{t})}$, and $\mathbf{K}(\mathbf{X}_{t},\mathbf{X}_{t})=[k(\bm{x},\widetilde{\bm{x}})]_{\bm{x},\widetilde{\bm{x}}\in\mathcal{E}(\mathbf{X}_{t})}$ are the kernel vector at $\bm{x}$, output vector up to step $t$, and the gram matrix, respectively. In the above definition, we denote $\mathcal{E}(\mathbf{X}_{t})$ as the subset of $\mathbf{X}_{t}$ such that the completely correlated inputs with one of the past inputs are eliminated. Namely, we define $\mathcal{E}(\mathbf{X}_{t})$ inductively as $\mathcal{E}(\mathbf{X}_{t})=\mathcal{E}(\mathbf{X}_{t-1})\cup\{\bm{x}_{t}\}$ if $\sigma^{2}(\bm{x}_{t};\mathbf{X}_{t-1})>0$; otherwise, $\mathcal{E}(\mathbf{X}_{t})=\mathcal{E}(\mathbf{X}_{t-1})$. Here, we define $\mathcal{E}(\mathbf{X}_{1})=\mathbf{X}_{1}$. Note that if there are no duplications in the input sequence $\bm{x}_{1},\ldots,\bm{x}_{t}$, then $\mathcal{E}(\mathbf{X}_{t})=\mathbf{X}_{t}$ holds under $k=k_{\mathrm{SE}}$ or $k=k_{\mathrm{Mat\acute{e}rn}}$. Furthermore, for the ease of notation, we set $\mu(\bm{x};\mathbf{X})=0$ and $\sigma^{2}(\bm{x};\mathbf{X})=k(\bm{x},\bm{x})$ for $\mathbf{X}=\emptyset$.
Maximum Information Gain.
Let us define the kernel-dependent complexity parameter $\gamma_{T}(\lambda^{2})$ as
@@
where $\lambda>0$ and $\bm{I}_{T}$ are any positive parameter and $T\times T$-identity matrix, respectively. The quantity $\gamma_{T}(\lambda^{2})$ is called maximum information gain (MIG) (Srinivas et al., 2010) since the quantity $\frac{1}{2}\ln\det(\bm{I}_{T}+\lambda^{-2}\mathbf{K}(\mathbf{X}_{T},\mathbf{X}_{T}))$ represents the mutual information between the underlying function $f$ and training outputs under the noisy-GP model with variance parameter $\lambda^{2}$. MIG plays an important role in the theoretical analysis of GP-bandits, and their increasing speed is analyzed in several commonly used kernels. For example, $\gamma_{T}(\lambda^{2})=O(\ln^{d+1}(T/\lambda^{2}))$ and $\gamma_{T}(\lambda^{2})=O((T/\lambda^{2})^{\frac{d}{2\nu+d}}\ln^{\frac{2\nu}{2\nu+d}}(T/\lambda^{2}))$ under $k=k_{\mathrm{SE}}$ and $k=k_{\mathrm{Mat\acute{e}rn}}$ with $\nu>1/2$, respectively [^1] These orders hold as $T\rightarrow\infty,\lambda\rightarrow 0$. (Vakili et al., 2021b).
@@
GP-upper confidence bound (GP-UCB) proposed in (Srinivas et al., 2010) is a well-known GP-bandit algorithm for noisy settings. Regarding the noise-free setting, Lyu et al. (2019) propose the noise-free version of GP-UCB, described in Algorithm 1.
@@
0:  Compact input domain $\mathcal{X}\subset\mathbb{R}^{d}$, RKHS norm upper bound $B\in(0,\infty)$.
1:  $\mathbf{X}_{0}\leftarrow\emptyset$, $\beta^{1/2}\leftarrow B$.
@@
4:     Observe $f(\bm{x}_{t})$ and update $\mathbf{X}_{t}$ and $\bm{f}_{t}(\mathbf{X}_{t})$.

### v2

@@ Abstract
We study the noise-free Gaussian Process (GP) bandit problem, in which a learner seeks to minimize regret through noise-free observations of a black-box objective function that lies in a known reproducing kernel Hilbert space (RKHS). The Gaussian Process Upper Confidence Bound (GP-UCB) algorithm is a well-known approach for GP bandits, where query points are adaptively selected based on the GP-based upper confidence bound score. While several existing works have reported the practical success of GP-UCB, its theoretical performance remains suboptimal. However, GP-UCB often empirically outperforms other nearly-optimal noise-free algorithms that use non-adaptive sampling schemes. This paper resolves the gap between theoretical and empirical performance by establishing a nearly-optimal regret upper bound for noise-free GP-UCB. Specifically, our analysis provides the first constant cumulative regret bounds in the noise-free setting for both the squared exponential kernel and the Matérn kernel with some degree of smoothness.
@@ 1 Introduction
This paper studies the noise-free Gaussian Process (GP) bandit problem, where the learner seeks to minimize regret through noise-free observations of the black-box objective function. Several existing works tackle this problem, and some of them (iwazaki2025improvedregretanalysisgaussian; salgiarandom) propose algorithms whose regret nearly matches the lower bound of (liregret). For ease of theoretical analysis, these algorithms rely on the non-adaptive sampling scheme, whose query points are chosen independently of the observed function values, such as the uniform sampling (salgiarandom) or maximum variance reduction (iwazaki2025improvedregretanalysisgaussian). Although the theoretical superiority of such non-adaptive algorithms is shown, in existing noisy setting literature (bogunovic2022robust; iwazakinear; li2022gaussian), their empirical performance has been reported to be worse than that of fully adaptive strategies such as GP upper confidence bound (GP-UCB) (srinivas10gaussian). Unsurprisingly, we also observe such empirical and practical gaps in a noise-free setting, as shown in Figure 1. These observations suggest the possibility of further theoretical improvement in the practical fully adaptive algorithm. From this motivation, our work aims to establish the nearly-optimal regret for GP-UCB, which is one of the well-known adaptive GP bandit algorithms, and its existing guarantees only show strictly sub-optimal regret in a noise-free setting (kim2024bayesian; lyu2019efficient).
@@
  • We give a refined regret analysis of GP-UCB (Theorems 1 and 2), which matches both the cumulative regret lower bounds of (liregret) and the simple regret lower bound of (bull2011convergence) up to polylogarithmic factors in the Matérn kernel. Regarding cumulative regret, our analysis shows that GP-UCB achieves the constant $O(1)$ regret under the squared exponential and Matérn kernel with $d>\nu$. The results are summarized in Tables 1 and 2.
  • Our key theoretical contribution is the new algorithm-independent upper bounds for the observed posterior standard deviations (Lemmas 35) by bridging the information gain-based analysis in the noisy regime to the noise-free setting. Furthermore, as discussed in Section 3, these results have the potential to translate existing confidence bound-based algorithms for noisy settings into nearly-optimal noise-free variants beyond the analysis of GP-UCB.
@@
Various existing works study the theory for the noisy GP bandits (chowdhury2017kernelized; li2022gaussian; scarlett2017lower; srinivas10gaussian; valko2013finite). Regarding noise-free settings, to our knowledge, (bull2011convergence) is the first work that shows both the upper bound and the lower bound for simple regret via the expected improvement (EI) strategy. After that, the analysis of the cumulative regret is shown in (lyu2019efficient) with GP-UCB. Recently, vakili2022open conjectures the lower bound of the cumulative regret in the noise-free setting, suggesting a superior algorithm exists in the noise-free setting. Later, the following work (liregret) formally validates the conjectured lower bound of (vakili2022open)  [^1] Although liregret considers the cascading structure of the observation process, the standard noise-free setting is the special case of their setting.. Motivated by the lower bounds, several works (flynn2024tighter; kim2024bayesian; iwazaki2025improvedregretanalysisgaussian; salgiarandom) studied the improved algorithm to achieve superior regret to the result of (lyu2019efficient). Although some of them propose the nearly-optimal algorithms (iwazaki2025improvedregretanalysisgaussian; salgiarandom), their algorithms are based on a non-adaptive sampling scheme, whose inferior performance has been reported in the existing work (bogunovic2022robust; iwazakinear; li2022gaussian). Here, our motivation is to establish the nearly-optimal theory under the fully adaptive nature of GP-UCB; however, our proof relates to the analysis in (iwazaki2025improvedregretanalysisgaussian) for the non-adaptive maximum variance reduction algorithm. Their analysis includes the noise-free setting as a special case and provides nearly-optimal regret under broader varying noise variance settings. However, its applicability is limited to the maximum variance reduction algorithm. Our analysis can be interpreted as a refined version of that in (iwazaki2025improvedregretanalysisgaussian) by focusing on the noise-free setting. Finally, although our paper studies the frequentist assumption that the underlying function is fixed, our core results (Lemmas 35) are also applicable to the regret analysis in the Bayesian setting, whose underlying function is drawn from a known GP (freitas2012exponential; grunewalder2010regret; Russo2014-learning; russo2014learning; scarlett2018tight; srinivas10gaussian).
[…]
Algorithm Regret (SE) Regret (Matérn) Type
@@
Kernel-AMM-UCB $O\left(\ln^{d+1}T\right)$ $\widetilde{O}\Bigl(T^{\frac{\nu d+d^{2}}{2\nu^{2}+2\nu d+d^{2}}}\Bigr)$ D
REDS N/A $\widetilde{O}\left(T^{\frac{d-\nu}{d}}\right)$ $O\left(\ln^{\frac{5}{2}}T\right)$ $O\left(\ln^{\frac{3}{2}}T\right)$ P
PE $O\left(\ln T\right)$ $\widetilde{O}\left(T^{\frac{d-\nu}{d}}\right)$ $O\left(\ln^{2+\alpha}T\right)$ $O\left(\ln T\right)$ D
@@
Lower Bound N/A $\Omega\left(T^{\frac{d-\nu}{d}}\right)$ $\Omega(1)$ $\Omega(1)$ N/A
Table 2: Comparison between existing noiseless algorithms’ guarantees for simple regret and our result (adapted from (iwazaki2025improvedregretanalysisgaussian)). In the regrets of GP-UCB+, EXPLOIT+, MVR, and GP-UCB, $\alpha>0$ and $C>0$ are arbitrary fixed constants and some positive constants, respectively.
Algorithm Regret (SE) Regret (Matérn) Type
@@
Kernel-AMM-UCB $O\left(\frac{\ln^{d+1}T}{T}\right)$ $\widetilde{O}\Bigl(T^{-\frac{\nu d+2\nu^{2}}{2\nu^{2}+2\nu d+d^{2}}}\Bigr)$ D
GP-UCB+, EXPLOIT+ $O\left(\exp\left(-CT^{\frac{1}{d}-\alpha}\right)\right)$ $O\left(T^{-\frac{\nu}{d}+\alpha}\right)$ P
MVR $O\left(\exp\left(-\frac{1}{2}T^{\frac{1}{d+1}}\ln^{-\alpha}T\right)\right)$ $\widetilde{O}\left(T^{-\frac{\nu}{d}}\right)$ D
GP-UCB $O\left(\sqrt{T}\exp\left(-\frac{1}{2}CT^{\frac{1}{d+1}}\right)\right)$ $\widetilde{O}\left(T^{-\frac{\nu}{d}}\right)$ D
(Our analysis)
@@
Figure 1: Empirical performance comparison of GP-UCB and two existing nearly-optimal algorithms: random exploration with domain shrinking (REDS) (salgiarandom) and phased elimination (PE) (iwazaki2025improvedregretanalysisgaussian). From left to right, the plots show the average cumulative regret over $3000$ independent runs under the squared exponential kernel, Matérn kernel ($\nu=5/2$), and Matérn kernel ($\nu=3/2$) with $d=2$, respectively. Here, $d$ and $\nu$ represent the dimension of the input and the smoothness parameter, respectively. Detailed experimental settings are provided in Appendix B.
@@ 2 Preliminaries
Noise-free GP bandit problem.
We study the GP bandit problem under noise-free observations. Let $\mathcal{X}\subset\mathbb{R}^{d}$ be a compact input domain, and consider a black-box objective function $f:\mathcal{X}\to\mathbb{R}$ that can only be evaluated point-wise. At each step $t\in\mathbb{N}_{+}$, the learner selects a query point $\bm{x}_{t}\in\mathcal{X}$ and observes its function value $f(\bm{x}_{t})$. After $T$ steps, the performance of the learner is measured using either the cumulative regret $R_{T}$ or the simple regret $r_{T}$, which are respectively defined as follows:
$R_{T}$ $=\sum_{t=1}^{T}f(\bm{x}^{\ast})-f(\bm{x}_{t}),$ (1)
@@
Here, $\bm{x}^{\ast}\in\mathrm{arg\,max}_{\bm{x}\in\mathcal{X}}f(\bm{x})$ is a global maximizer, and $\widehat{\bm{x}}_{T}$ denotes the estimated maximizer returned by the algorithm at the end of the final step $T$.
@@
To construct the GP-bandit algorithm, the GP model (Rasmussen2005-Gaussian) plays a central role in balancing the trade-off between exploration and exploitation. First, we adopt a Gaussian process prior with zero mean and covariance (kernel) function $k:\mathcal{X}\times\mathcal{X}\to\mathbb{R}$. Then, given a sequence of queried points $\mathbf{X}_{t}=(\bm{x}_{1},\dots,\bm{x}_{t})$ and their corresponding evaluations $\bm{f}(\mathbf{X}_{t})=(f(\bm{x}_{1}),\dots,f(\bm{x}_{t}))$, the posterior mean $\mu(\bm{x};\mathbf{X}_{t})$ and variance $\sigma^{2}(\bm{x};\mathbf{X}_{t})$ of $f(\bm{x})$ at a new point $\bm{x}\in\mathcal{X}$ are
@@
where $\bm{k}(\bm{x},\mathcal{E}(\mathbf{X}_{t}))=[k(\widetilde{\bm{x}},\bm{x})]_{\widetilde{\bm{x}}\in\mathcal{E}(\mathbf{X}_{t})}$ is the kernel vector, $\mathbf{K}(\mathcal{E}(\mathbf{X}_{t}),\mathcal{E}(\mathbf{X}_{t}))$ is the Gram matrix, and $\bm{f}(\mathcal{E}(\mathbf{X}_{t}))=[f(\widetilde{\bm{x}})]_{\widetilde{\bm{x}}\in\mathcal{E}(\mathbf{X}_{t})}$. In the above definition, $\mathcal{E}(\mathbf{X}_{t})$ denotes the subset of $\mathbf{X}_{t}$ obtained by removing any fully correlated inputs (i.e., with zero posterior variance) with previous ones. Namely, we define $\mathcal{E}(\mathbf{X}_{t})$ inductively as $\mathcal{E}(\mathbf{X}_{t})=\mathcal{E}(\mathbf{X}_{t-1})\cup\{\bm{x}_{t}\}$ if $\sigma^{2}(\bm{x}_{t};\mathbf{X}_{t-1})>0$; otherwise, $\mathcal{E}(\mathbf{X}_{t})=\mathcal{E}(\mathbf{X}_{t-1})$. Here, we define $\mathcal{E}(\mathbf{X}_{1})=\mathbf{X}_{1}$. Note that if there are no duplications in the input sequence $\bm{x}_{1},\ldots,\bm{x}_{t}$, then $\mathcal{E}(\mathbf{X}_{t})=\mathbf{X}_{t}$ holds under commonly used kernel (covariance) functions, such as squared exponential and Matérn kernels (precisely defined in the next paragraph). Furthermore, for the ease of notation, we set $\mu(\bm{x};\mathbf{X})=0$ and $\sigma^{2}(\bm{x};\mathbf{X})=k(\bm{x},\bm{x})$ for $\mathbf{X}=\emptyset$.
Kernel function and information gain.
Regarding the choice of the kernel function, we focus on the squared exponential (SE) kernel
[…]
and the Matérn kernel
[…]
where $\ell>0$ and $\nu>0$ are the lengthscale and smoothness parameter, respectively. Furthermore, $J_{\nu}$ and $\Gamma$ are the modified Bessel and Gamma functions, respectively. These two kernels are commonly used and analyzed in GP-bandits (scarlett2017lower; srinivas10gaussian). The convergence rate of the function estimation in GP regression depends on the choice of kernel, which in turn affects the resulting regret upper bound. Therefore, to capture the problem complexity in a kernel-dependent manner, the following kernel-dependent information theoretic quantity $\gamma_{T}(\lambda^{2})$ is often employed in the analysis of GP bandits:
@@
where $\lambda>0$ and $\bm{I}_{T}$ are any positive parameter and $T\times T$-identity matrix, respectively. The quantity $\gamma_{T}(\lambda^{2})$ is called maximum information gain (MIG) (srinivas10gaussian) since the quantity $\frac{1}{2}\ln\det(\bm{I}_{T}+\lambda^{-2}\mathbf{K}(\mathbf{X}_{T},\mathbf{X}_{T}))$ represents the mutual information between the underlying function $f$ and training outputs under the noisy-GP model with variance parameter $\lambda^{2}$. The increasing speed of MIG is analyzed in several commonly used kernels. For example, $\gamma_{T}(\lambda^{2})=O(\ln^{d+1}(T/\lambda^{2}))$ and $\gamma_{T}(\lambda^{2})=\widetilde{O}((T/\lambda^{2})^{\frac{d}{2\nu+d}})$ under $k=k_{\mathrm{SE}}$ and $k=k_{\mathrm{Mat\acute{e}rn}}$ with $\nu>1/2$, respectively (vakili2021information) [^2] These orders hold as $T\rightarrow\infty,\lambda\rightarrow 0$..
Regularity assumption.
It is hopeless to derive meaningful guarantees without further assumptions about $f$. To obtain a valid estimate of $f$ through the GP-model, we assume that $f$ lies in the known reproducing kernel Hilbert space (RKHS) (aronszajn1950theory) corresponding to the kernel $k$ and has bounded norm. The formal description is given below.
Assumption 1.
The objective function $f$ lies in the reproducing kernel Hilbert space (RKHS) associated with a known positive definite kernel $k:\mathcal{X}\times\mathcal{X}\rightarrow\mathbb{R}$. We assume that $k(\bm{x},\bm{x})\leq 1$ for all $\bm{x}\in\mathcal{X}$, and that the RKHS norm $\|f\|_{k}$ satisfies $\|f\|_{k}\leq B<\infty$.
This is a standard assumption in the GP-bandit literature (chowdhury2017kernelized; scarlett2017lower; srinivas10gaussian), and is leveraged to derive the confidence bound of $f$ (kanagawa2018gaussian; lyu2019efficient; vakili2021optimal). Here, RKHS associated with kernel $k$ is given as the closure of the linear span: $\mathcal{H}_{k}^{(\mathrm{pre})}\coloneqq\{\sum_{i=1}^{n}c_{i}k(\bm{x}^{(i)},\cdot)\mid n\in\mathbb{N}_{+},c_{1},\ldots,c_{n}\in\mathbb{R},\bm{x}^{(1)},\ldots,\bm{x}^{(n)}\in\mathcal{X}\}$ (kanagawa2018gaussian). Therefore, intuitively, under Assumption 1, we can interpret that the basic properties of the objective function $f$, such as continuity and smoothness, are encoded through the choice of the kernel function $k$.
@@
GP-UCB (srinivas10gaussian) is a widely used algorithm for the noisy GP bandit setting. Its noiseless variant was proposed by lyu2019efficient, and is outlined in Algorithm 1. Their analysis largely follows the framework developed for the noisy case in (srinivas10gaussian), but differs in the confidence bound, which is strictly tighter than that in the noisy setting. Although this refinement of the confidence bound leads to superior regret compared with the noisy setting, the resulting regret is still strictly sub-optimal in the noise-free setting. This fact suggests that we require the other fundamental modification from the proof in (srinivas10gaussian).
@@
0: Compact input domain $\mathcal{X}\subset\mathbb{R}^{d}$, Kernel function $k$, and RKHS norm upper bound $B\in(0,\infty)$.
[…]
@@
4:  Observe $f(\bm{x}_{t})$ and update the posterior mean and variance.